# Scalable Sobolev IPM for Probability Measures on a Graph

### v1

@@ Appendix A Proofs
$\hat{{\mathcal{S}}}_{p}(\mu,\nu)$ $=\sup_{g\in L^{p^{\prime}}({\mathbb{G}},\lambda):\,\left\lVert g\right\rVert_{L^{p^{\prime}}_{\hat{w}}}\leq 1}\left|\int_{{\mathbb{G}}}\hat{w}(x)\hat{f}(x)g(x)\lambda(\mathrm{d}x)\right|$
@@ C.1 Further Results
We follow the same reasoning for the proof of Theorem 3.4. We recall Equation (LABEL:app:eq:dualnorm_regSobolevIPM)
@@ C.2 Further Experimental Results
  • Figures 7 and 8 illustrate SVM results and time consumption for kernel matrices on graphs ${\mathbb{G}}_{\text{Sqrt}}$ and ${\mathbb{G}}_{\text{Log}}$ respectively where the number of nodes $M=1000$.
  • Figures 9 and 10 illustrate SVM results and time consumption for kernel matrices on graphs ${\mathbb{G}}_{\text{Sqrt}}$ and ${\mathbb{G}}_{\text{Log}}$ respectively where the number of nodes $M=100$.
@@
  • Figures 11 and 12 illustrate SVM results and time consumption for kernel matrices on graphs ${\mathbb{G}}_{\text{Sqrt}}$ and ${\mathbb{G}}_{\text{Log}}$ respectively where the number of nodes $M=1000$.
  • Figures 13 and 14 illustrate SVM results and time consumption for kernel matrices on graphs ${\mathbb{G}}_{\text{Sqrt}}$ and ${\mathbb{G}}_{\text{Log}}$ respectively where the number of nodes $M=100$.
@@
  • Figures 16 and 17 illustrate SVM results and time consumption for kernel matrices on graphs ${\mathbb{G}}_{\text{Sqrt}}$ and ${\mathbb{G}}_{\text{Log}}$ respectively where the number of nodes $M=1000$.
  • Figures 18 and 19 illustrate SVM results and time consumption for kernel matrices on graphs ${\mathbb{G}}_{\text{Sqrt}}$ and ${\mathbb{G}}_{\text{Log}}$ respectively where the number of nodes $M=100$.
@@
  • Figures 21(a) and 22(a) illustrate SVM results and time consumption for kernel matrices on graphs ${\mathbb{G}}_{\text{Sqrt}}$ and ${\mathbb{G}}_{\text{Log}}$ respectively where the number of nodes $M=1000$.
  • Figures 21(b) and 22(b) illustrate SVM results and time consumption for kernel matrices on graphs ${\mathbb{G}}_{\text{Sqrt}}$ and ${\mathbb{G}}_{\text{Log}}$ respectively where the number of nodes $M=100$.
Figure 7: SVM results and time consumption for kernel matrices in document classification with graph ${\mathbb{G}}_{\text{Sqrt}}$ where the number of nodes $M=1000$.
Figure 8: SVM results and time consumption for kernel matrices in document classification with graph ${\mathbb{G}}_{\text{Log}}$ where the number of nodes $M=1000$.
Figure 9: SVM results and time consumption for kernel matrices in document classification with graph ${\mathbb{G}}_{\text{Sqrt}}$ where the number of nodes $M=100$.
Figure 10: SVM results and time consumption for kernel matrices in document classification with graph ${\mathbb{G}}_{\text{Log}}$ where the number of nodes $M=100$.
Figure 11: SVM results and time consumption for kernel matrices in TDA with graph ${\mathbb{G}}_{\text{Sqrt}}$ where the number of nodes $M=1000$.
Figure 12: SVM results and time consumption for kernel matrices in TDA with graph ${\mathbb{G}}_{\text{Log}}$ where the number of nodes $M=1000$.
Figure 13: SVM results and time consumption for kernel matrices in TDA with graph ${\mathbb{G}}_{\text{Sqrt}}$ where the number of nodes $M=100$.
Figure 14: SVM results and time consumption for kernel matrices in TDA with graph ${\mathbb{G}}_{\text{Log}}$ where the number of nodes $M=100$.
Figure 15: SVM results and time consumption for kernel matrices of slice variants with graph ${\mathbb{G}}_{\text{Sqrt}}$ where the number of nodes $M=10^{4}$.
Figure 16: SVM results and time consumption for kernel matrices of slice variants with graph ${\mathbb{G}}_{\text{Sqrt}}$ where the number of nodes $M=1000$.
Figure 17: SVM results and time consumption for kernel matrices of slice variants with graph ${\mathbb{G}}_{\text{Log}}$ where the number of nodes $M=1000$.
Figure 18: SVM results and time consumption for kernel matrices of slice variants with graph ${\mathbb{G}}_{\text{Sqrt}}$ where the number of nodes $M=100$.
Figure 19: SVM results and time consumption for kernel matrices of slice variants with graph ${\mathbb{G}}_{\text{Log}}$ where the number of nodes $M=100$.
@@ C.3 Further Discussions
$\bullet$ Compute distance between two (different) input graphs. For examples, Petric Maretic et al. (2019); Dong & Sawin (2020); Tam & Dunson (2022) compute OT problem (i.e., a specific instance of IPM) between two (different) input graphs, where their goals are to compute distance between two input graphs. They are essentially different to our considered problem which computes distance (i.e., regularized Sobolev IPM) between two input probability measures supported on the same graph.
@@
The number of pairs in training and test for kernel SVM.

### v2

@@ 7 Conclusion
Acknowledgements
We thank the area chairs and anonymous reviewers for their comments. KF has been supported in part by Grant-in-Aid for Transformative Research Areas (A) 22H05106 and JST CREST JPMJCR2015. HH acknowledges the support of JSPS KAKENHI JP25H01494 and JP23K24909. TL gratefully acknowledges the support of JSPS KAKENHI Grant number 23K11243, and Mitsui Knowledge Industry Co., Ltd. grant.
@@ Appendix A Proofs
$\hat{{\mathcal{S}}}_{p}(\mu,\nu)$ $=\sup_{g\in L^{p^{\prime}}({\mathbb{G}},\lambda):\,\left\lVert g\right\rVert_{L^{p^{\prime}}_{\hat{w}}}\leq 1}\left|\int_{{\mathbb{G}}}\hat{w}(x)\hat{f}(x)g(x)\lambda(\mathrm{d}x)\right|$ (23)
@@ B.3 A Review on IPM and Wasserstein Distance
$\bullet$ Tree topology.
When graph ${\mathbb{G}}$ is a tree, the $1$-Wasserstein distance yields a closed-form expression for a fast computation (Le et al., 2019) (i.e., tree-Wasserstein (TW)). Le & Nguyen (2021) extend it for unbalanced input measures (i.e., which may have different total masses), and propose a novel regularization approach, named regularized entropy partial transport which also admits a closed-form formula for a fast computation. Furthermore, Tran et al. (2025a, b, c, d) have recently developed tree-sliced-Wasserstein on tree-structured systems of lines (i.e., tree systems) to extend TW for applications with dynamic-support measures including generative models, and diffusion models.
@@ C.1 Further Results
We follow the same reasoning for the proof of Theorem 3.4. We recall Equation (23)
@@ C.2 Further Experimental Results
  • Figures 8 and 9 illustrate SVM results and time consumption for kernel matrices on graphs ${\mathbb{G}}_{\text{Sqrt}}$ and ${\mathbb{G}}_{\text{Log}}$ respectively where the number of nodes $M=1000$.
  • Figures 10 and 11 illustrate SVM results and time consumption for kernel matrices on graphs ${\mathbb{G}}_{\text{Sqrt}}$ and ${\mathbb{G}}_{\text{Log}}$ respectively where the number of nodes $M=100$.
@@
  • Figures 12 and 13 illustrate SVM results and time consumption for kernel matrices on graphs ${\mathbb{G}}_{\text{Sqrt}}$ and ${\mathbb{G}}_{\text{Log}}$ respectively where the number of nodes $M=1000$.
  • Figures 14 and 15 illustrate SVM results and time consumption for kernel matrices on graphs ${\mathbb{G}}_{\text{Sqrt}}$ and ${\mathbb{G}}_{\text{Log}}$ respectively where the number of nodes $M=100$.
@@
  • Figures 17 and 18 illustrate SVM results and time consumption for kernel matrices on graphs ${\mathbb{G}}_{\text{Sqrt}}$ and ${\mathbb{G}}_{\text{Log}}$ respectively where the number of nodes $M=1000$.
  • Figures 19 and 20 illustrate SVM results and time consumption for kernel matrices on graphs ${\mathbb{G}}_{\text{Sqrt}}$ and ${\mathbb{G}}_{\text{Log}}$ respectively where the number of nodes $M=100$.
@@
  • Figures 22(a) and 23(a) illustrate SVM results and time consumption for kernel matrices on graphs ${\mathbb{G}}_{\text{Sqrt}}$ and ${\mathbb{G}}_{\text{Log}}$ respectively where the number of nodes $M=1000$.
  • Figures 22(b) and 23(b) illustrate SVM results and time consumption for kernel matrices on graphs ${\mathbb{G}}_{\text{Sqrt}}$ and ${\mathbb{G}}_{\text{Log}}$ respectively where the number of nodes $M=100$.
Figure 8: SVM results and time consumption for kernel matrices in document classification with ${\mathbb{G}}_{\text{Sqrt}}$ and $M=1000$.
Figure 9: SVM results and time consumption for kernel matrices in document classification with ${\mathbb{G}}_{\text{Log}}$ and $M=1000$.
Figure 10: SVM results and time consumption for kernel matrices in document classification with ${\mathbb{G}}_{\text{Sqrt}}$ and $M=100$.
Figure 11: SVM results and time consumption for kernel matrices in document classification with ${\mathbb{G}}_{\text{Log}}$ and $M=100$.
Figure 12: SVM results and time consumption for kernel matrices in TDA with ${\mathbb{G}}_{\text{Sqrt}}$ and $M=1000$.
Figure 13: SVM results and time consumption for kernel matrices in TDA with ${\mathbb{G}}_{\text{Log}}$ and $M=1000$.
Figure 14: SVM results and time consumption for kernel matrices in TDA with ${\mathbb{G}}_{\text{Sqrt}}$ and $M=100$.
Figure 15: SVM results and time consumption for kernel matrices in TDA with ${\mathbb{G}}_{\text{Log}}$ and $M=100$.
Figure 16: SVM results and time consumption for kernel matrices of slice variants with ${\mathbb{G}}_{\text{Sqrt}}$ and $M=10^{4}$.
Figure 17: SVM results and time consumption for kernel matrices of slice variants with ${\mathbb{G}}_{\text{Sqrt}}$ and $M=1000$.
Figure 18: SVM results and time consumption for kernel matrices of slice variants with ${\mathbb{G}}_{\text{Log}}$ and $M=1000$.
Figure 19: SVM results and time consumption for kernel matrices of slice variants with ${\mathbb{G}}_{\text{Sqrt}}$ and $M=100$.
Figure 20: SVM results and time consumption for kernel matrices of slice variants with ${\mathbb{G}}_{\text{Log}}$ and $M=100$.
@@ C.3 Further Discussions
$\bullet$ Compute distance between two (different) input graphs. For examples, Petric Maretic et al. (2019); Dong & Sawin (2020); Le et al. (2021); Tam & Dunson (2022) compute OT problem (i.e., a specific instance of IPM) between two (different) input graphs, where their goals are to compute distance between two input graphs. They are essentially different to our considered problem which computes distance (i.e., regularized Sobolev IPM) between two input probability measures supported on the same graph.
@@
The number of pairs in training and test for kernel SVM (Le et al., 2024).
@@
Distinguish to the approach in Peyre (2018).
We study the Sobolev IPM for probability measures supported on a given graph. The Sobolev IPM (Equation (4)) constrains a critic function within the unit ball defined by the Sobolev norm (Equation (3)) involving both the critic function and its gradient, while Peyre (2018) considers the weighted homogeneous Sobolev norm which constraints a critic function in a semi-norm involving only gradient of the critic function Peyre (2018, Equations (3) and (4))Peyre (2018)’s approach shares the same spirit as the $2$-order Sobolev transport (Le et al., 2022).
The proposed regularized Sobolev IPM may share some spirit with the weighted homogeneous Sobolev norm and the Sobolev transport (i.e., constraint on gradient of a critic function). However, we clarify that the weight function (Equation (5)) plays the key role to establish the equivalence between Sobolev norm and weighted norm (Theorem 3.2) for our novel regularization for Sobolev IPM.
Further discussions on the research problem and approach.
We emphasize that our approach aims neither to derive a tight bound for the gap nor to provide a sharp approximation for the challenging optimization problem (Equation (4)) of the original Sobolev IPM. Our purpose is in a different direction. More precisely, we instead propose a novel regularization for Sobolev IPM which yields an equivalent metric to the original Sobolev IPM, and a closed-form expression for a fast computation. Note that, to our knowledge, there are no efficient algorithms to compute Sobolev IPM effectively. We believe that deriving tight bound or providing sharp approximation for the Sobolev IPM is interesting research direction, and is left for future investigation.